one-half of leo's age two years from now plus one-third of his age three years ago is Twenty-Five years.how old is he now?
Answer by Maths68(1474)

Let
Present age of Leo = x years old

Two years from now
Leo will be = x+2 years old
One-half of Leo's age then = (x+2)/2
Three Year's ago
Leo was = x-3 years old
One-third of Leo's age then = (x-3)/3

One-half of Leo's age from now + One-third of Leo's age three years ago = 25
(x+2)/2+(x-3)/3=25
Take LCD
(3(x+2)+2(x-3))/6=25
(3x+6+2x-6)/6=25
5x/6=25
5x=25*6
5x=150
x=150/5
x=30
Present age of Leo = x = 30 years old